Write the formula for an ionic compound formed from O2 with each of the following ions. magnesium ion Mgo sodium ion Na,O aluminum ion ALO,
Answer #1

Oxygen oxidation number -2, Magenisium oxidation number +2. So the compound magenisium oxide having formula MgO (+2-2=0).

Oxygen oxidation number -2, sodium oxidation number +1. So to balance the oxidation of oxygen two sodium atoms (2x+1= +2) are needed. So sodium oxide molecular formula Na2O (+2-2=0).

Aluminium oxidation number is +3 and oxygen oxidation number -2. So the molecular formula aluminium oxide Al2O3.

Al2 = +3 x 2 . = . + 6

O3 . = -2 x 3 = . -6

Al2O3 = +6-6 = 0.